# A backdoor attack against link prediction tasks with graph neural networks

### v1

@@ 5.2 The effects of different poisoning rates
Cora 0.2 93.66% 96.34% 90.88% 73.38% 0.11% -0.57% 0.45% -0.98%
0.5 98.40% 98.62% 96.66% 79.46% 0.52% -0.34% -0.01% 0.24%
1 99.53% 99.58% 97.66% 89.40% 0.02% -0.13% -0.33% -0.19%
2 99.06% 99.18% 98.26% 88.58% 0.27% 0.76% -0.19% 0.78%
5 99.62% 99.92% 99.16% 99.16% -0.18% 0.13% 0.07% 1.63%
CiteSeer 0.2 99.06% 98.94% 99.12% 96.48% 0.24% 0.46% -0.06% 0.43%
0.5 99.62% 99.90% 99.32% 98.72% -0.26% -0.70% -0.21% 0.19%
1 99.85% 99.84% 99.77% 99.63% 0.67% 0.64% 0.24% 1.08%
2 99.94% 99.98% 99.36% 99.58% 0.25% -0.15% 0.30% 0.08%
5 100.00% 99.98% 99.94% 99.93% 0.99% 0.12% 0.39% 0.11%
CS 0.2 95.08% 95.02% 96.08% 88.00% 0.56% 0.14% 0.02% -0.16%
0.5 97.48% 96.46% 97.56% 94.24% 0.04% -0.08% 0.02% -0.22%
1 97.89% 97.97% 98.22% 96.44% 0.04% -0.62% -0.06% 0.09%
2 98.62% 98.58% 98.64% 97.64% 0.15% 0.32% 0.12% 0.39%
5 99.10% 99.82% 98.92% 98.52% 0.17% 0.11% 0.06% -0.37%
Physics 0.2 98.78% 98.11% 93.95% 93.34% 0.36% -0.15% 0.07% 0.38%
0.5 99.32% 99.47% 97.73% 96.77% -0.61% 0.04% -0.04% -0.38%
1 99.59% 99.48% 98.62% 97.45% -0.25% 0.37% 0.02% -0.27%
2 99.79% 99.68% 99.76% 98.95% 0.25% 0.27% -0.11% 0.61%
5 100.00% 99.90% 100.00% 98.06% 0.22% 0.16% 0.17% 0.03%
@@
Figure 3 is a visualization of the above experimental results, in which the vertical coordinates of the top four plots are the ASR, the vertical coordinates of the bottom four plots are the BPD, and the horizontal coordinates are the poisoning rates. From Figure 3, we can observe that as p increases, the ASRs of the backdoor models increases over the four datasets. Taking the Cora dataset as an example, when p=0.2%, the attack success rates of the four backdoor models GAE, VGAE, ARGA and ARVGA are 93.66%, 96.34%, 90.88%,73.38% respectively, and their ASRs are increased to 99.53%, 99.58%, 97.66%,89.40% respectively when p=2%. Thus, our method requires only a small poisoning rate to achieve a good attack effect, which ensures that our attack is not easily detected.
From Figure 3, we can also observe that the BPD of the four backdoor models in the four datasets fluctuates somewhat with $p$, but is overall less than 1%. That indicates that the backdoor models generated by the attack method proposed in this paper have prediction accuracies for clean samples that are not much different from those of the clean models, which makes the backdoor models very stealthy.
5.3 Comparison with baseline
In this section we compare our backdoor attack method with baseline. We compare the attack performance of the backdoor models against baseline for the four link prediction models (GAE, VGAE, ARGA and ARVGA) on four datasets (Cora, CiteSeer, CS and Physcis), and then compare the attack performance of the backdoor model against Baseline for the GAE model on the large-scale dataset physics .
1.Comparison of attack performance against baseline on three datasets (Cora, CiteSeer, and CS).
We test the performance of our backdoor attack on three datasets (Cora, CiteSeer, and CS) for four link prediction models (GAE, VGAE, ARGA and ARVGA) and compared it with the experimental results of BASELINE as shown in table 4, where, the experimental results of baseline come from the original paper [22], whose experiments were conducted under the condition of 10% poisoning rate. Since our method requires only small poisoning rates to achieve high-precision attack effectiveness, we choose the experimental results with the 1% poisoning rate in table 4 to compare with the baseline.
Table 4: Comparison with the baseline.
Datasets models ASR BPD
@@ 5.3 Comparison with baseline
Cora GAE 81.25% 99.53% 2.82% 0.02%
VGAE 83.22% 99.58% 4.65% -0.13%
ARGA 92.56% 97.66% 1.08% -0.33%
ARVGA 85.25% 89.40% 2.49% -0.19%
CiteSeer GAE 82.25% 99.85% 3.13% 0.67%
VGAE 89.59% 99.84% 8.51% 0.64%
ARGA 98.78% 99.77% 4.97% 0.24%
ARVGA 81.55% 99.63% 1.60% 1.08%
CS GAE 50.61% 97.89% 1.13% 0.04%
VGAE 67.26% 97.97% 0.95% -0.62%
ARGA 63.52% 98.22% 0.32% -0.06%
ARVGA 68.67% 96.44% 2.86% 0.09%
From the experimental results in Table 4, the backdoor attack method proposed in this paper against the four target models outperforms baseline on all three datasets, achieving more than 89% ASRs on all of them. Taking the Cora dataset as an example, our method achieved 99.53%, 99.58%, 97.66%, and 89.40% ASRs on the four link prediction models based on GAE, VGAE, ARGA and ARVGA, respectively, which are higher than the baseline’s 81.25%, 83.22%, 92.56%, and 85.25%, respectively. All of our results on the Citeseer dataset achieve an ASR of more than 99%. On the CS dataset, the ASRs of baseline drops dramatically, while our method still achieves an over 96% ASRs at CS.
2. Comparison of backdoor attack performance with Baseline on the large-scale dataset
Considering that graph data in the real world are complex, Baseline specifically evaluates the effectiveness and evasiveness on the large-scale dataset physcics. Since baseline only tested the backdoor model for the GAE model with the poisoning rate of 10%, we chose the data related to the GAE backdoor model with the poisoning rate of 1% on the physcics dataset from Table 3 for comparison with it, as shown in the table 5.
From Table 5, we can see that baseline’s attack performance on large-scale datasets is lower than ours, and our proposed backdoor attack method still has high ASRs and evasiveness on the large-scale dataset.
Table 5: Comparison with the baseline on the large-scale dataset
Method ASR BPD
baseline 39.02% 0.78%
ours 99.86% 0.13%
5.4 Testing the impact of different budgets on ASR and BPD
In this section we evaluate the impact of different budgets $k$ on ASR and BPD. We test the effect of different budgets on our backdoor attack on four datasets (Cora, CiteSeer, CS, and Physics) for four link predictions models (GAE, VGAE, ARGA and ARVGA). In the experiments on each dataset,$\lambda$ takes values of 0.2%, 0.5%, 1%, 2%, and 5%, respectively. The experimental results are shown in table 6.
Table 6: The impact of budgets on ASR and BPD.
Datasets $\lambda$ ASR BPD
@@ 5.4 Testing the impact of different budgets on ASR and BPD
Cora 0.2 96.28% 97.24% 82.57% 65.81% 0.47% -0.21% -0.04% 0.10%
0.5 99.21% 99.38% 92.41% 71.43% 0.07% -0.01% 0.17% 0.11%
1 99.53% 99.58% 97.66% 89.40% 0.02% -0.13% -0.33% -0.19%
2 99.68% 99.71% 99.04% 96.53% -0.62% $0$08% 0.46%
5 99.80% 99.89% 99.69% 99.14% 0.80% -0.17% -0.30% -0.85%
CiteSeer 0.2 99.38% 99.78% 97.17% 72.58% 0.72% -0.13% 0.21% 0.36%
0.5 99.95% 99.99% 99.53% 93.56% 0.96% 0.52% 0.01% 0.14%
1 99.85% 99.84% 99.77% 99.63% 0.67% 0.64% 0.24% 1.08%
2 99.95% 99.99% 99.85% 99.83% 0.02% -0.53% 0.70% -0.14%
5 100.00% 99.93% 99.89% 99.90% 0.60% 0.30% -0.78% -0.07%
CS 0.2 89.03% 87.76% 83.28% 70.86% -0.57% -0.18% $0$06%
0.5 94.96% 95.24% 95.39% 88.75% -0.20% 0.01% 0.12% -0.40%
1 97.89% 97.97% 98.22% 96.44% 0.04% -0.62% -0.06% 0.09%
2 98.10% 98.68% 99.02% 98.42% 0.14% 0.30% -0.01% 0.27%
5 98.53% 99.42% 99.49% 99.33% 0.73% -0.20% 0.01% 0.57%
Physics 0.2 95.89% 97.80% 86.21% 71.70% -0.56% -0.28% 0.02% 0.07%
0.5 98.74% 99.09% 96.31% 89.42% -0.72% -0.01% -0.16% 0.11%
1 99.59% 99.48% 98.62% 97.45% -0.25% 0.37% 0.02% -0.27%
2 100.00% 100.00% 99.53% 99.05% -0.33% -0.12% -0.05% -0.11%
5 99.92% 99.95% 99.87% 99.81% -0.41% 0.14% -0.06% 0.62%
Figure 4: The impact of budgets on ASR and BPD.
[…]
5.5 Ablation Studies
In order to verify the effect of the poisoning target node pairs selection method on the ASRs of the backdoor models,we conducted the following ablation studies. We generate poisoned training datasets by randomly selecting target node pairs to inject the trigger in the training sets of the four datasets (Cora, CiteSeer, CS and Physics), and retrain the four backdoor models and test their ASRs by using these poisoning training datasets respectively, and then compare the ASRs of the four backdoor models trained by selecting the target node pairs to inject the trigger through the NPS function as proposed in this paper, and the results of the experiments are shown in figeure 5. From the figure, we can see that the ASRs of random method is lower than that of ours. Especially on top of larger-scale graphs, such as CS and Physics, the ASRs of the Random method drop significantly, while our proposed NPS method still achieves excellent results. This is because our node pairs selection method chooses node pairs with sparse features, that are useful for our backdoor attack. The experimental results show that the poisoning method of injecting trigger by selecting target node pairs through the NPS function can effectively improve the ASRs of the backdoor attack.
Figure 5: Ablation studies.
6 Conclusion
@@
This paper discusses the feasibility of implementing the backdoor attack in link prediction. Specifically, we implement a single-node injection backdoor attack strategy that exploits the differences in the features distribution of the dataset to select the optimal set of features for the trigger node and utilizes the features of the node pairs to find the suitlable node pairs for poisoning. Our empirical evaluation results on four representative link prediction models (GAE, VGAE, ARGA and ARVGA) and four benchmark datasets show that our backdoor attack achieves high ASRs with little impact on the GNNs accuracies of clean test graphs. Currently there is little research on graph backdoor defense, and in the future we will investigate defense methods against link prediction backdoor attacks.

### v2

@@ V-B Testing the impact of different poisoning rates on ASR and BPD
Cora 0.2 95.39 94.59 91.55 74.82 -0.31 0.32 -0.64 -0.01
0.5 98.84 98.43 96.32 79.76 -0.64 0.28 -0.62 -0.16
1 99.00 99.13 97.53 88.16 -1.30 0.17 -0.71 -0.27
2 99.50 99.33 97.87 91.79 -1.24 -0.18 0.62 -0.62
5 99.76 99.61 98.76 94.72 -0.80 0.10 -0.62 -0.27
CiteSeer 0.2 98.45 98.47 98.41 95.85 0.30 0.65 -0.43 0.42
0.5 99.42 98.63 96.44 97.69 -0.28 -0.38 -0.40 0.19
1 99.69 99.85 99.81 98.88 0.32 0.02 -0.40 0.37
2 99.77 99.87 97.97 99.10 0.48 -0.34 -0.48 0.49
5 99.90 99.95 99.95 99.49 -0.78 -0.49 -0.24 0.54
CS 0.2 94.12 96.07 94.59 87.51 0.50 0.13 -0.02 0.25
0.5 97.38 98.03 96.85 91.00 0.76 0.08 -0.02 0.25
1 98.36 98.63 97.72 93.72 0.59 0.03 -0.03 0.30
2 99.77 98.87 99.97 99.10 0.46 -0.37 -0.48 0.49
5 99.91 99.95 99.95 99.49 -0.73 -0.50 -0.24 0.54
Physics 0.2 98.65 99.19 93.93 93.19 -0.44 -0.04 0.11 -0.65
0.5 99.81 98.96 95.86 96.28 -0.41 -0.01 0.11 -0.56
1 99.94 99.73 98.09 97.10 -0.44 -0.05 0.10 -0.59
2 99.95 99.78 99.02 98.12 -0.46 0.01 0.10 -0.56
5 99.97 99.88 99.31 98.51 -0.43 0.06 0.08 0.47
@@
Figure 3 is a visualization of the above experimental results, in which the vertical coordinates of the top four charts are the ASR, the vertical coordinates of the bottom four charts are the BPD, and the horizontal coordinates of all charts are the poisoning rates. From figure 3, we can observe that as p increases, the ASRs of the backdoored models increases over the four datasets. When the poisoning rate is increased to 2%, our attack achieves excellent attack performance on all datasets. Taking the Cora dataset as an example, when $p=0.2\%$, the ASRs of the four backdoored models GAE, VGAE, ARGA and ARVGA are 95.39%, 94.59%, 91.55%, 74.82% respectively, and their ASRs are increased to 99.50%, 99.33%, 97.87%, 91.79% respectively when $p=2\%$. Thus, the proposed backdoor attack can achieves good attack performance with a small poisoning rate, which ensures that our attack is effective and stealthy.
From figure 3, we can also observe that the BPD of the four backdoored models in the four datasets fluctuates somewhat with $p$, but the fluctuation is overall less than 1%. The results indicate that the backdoored models have close prediction accuracy to that of clean models on benign samples, which makes the backdoored models very stealthy.
V-C Comparison with baseline
In order to evaluate the proposed backdoor attack,we compare it with the only one state-of-the-art baseline LB[20] which is a backdoor attack method for link prediction tasks on static graphs ,using subgraphs as triggers. As the baseline has conducted backdoor attack performance tests on models(GAE, VGAE, ARGA and ARVGA) across three datasets(Cora, CiteSeer, CS), and also performed a backdoor attack performance test on model GAE on the large scale dataset Physics, we have compared the performance of our proposed backdoor attack method with the baseline accordingly.
1. Comparison of the performance of the proposed backdoor attack with that of the baseline on four models(GAE, VGAE, ARGA and ARVGA) across three datasets(Cora, CiteSeer, CS).
We test the performance of our backdoor attack on three datasets (Cora, CiteSeer, and CS) for four link prediction models (GAE, VGAE, ARGA and ARVGA) and compared it with the experimental results of baseline, where, the experimental results of baseline come from the original paper[20] and they were obtained with a poisoning rate of 10% poisoning rate. Since our method requires only small poisoning rates to achieve attack success rate, we choose the experimental results with the 1% poisoning rate to compare with the baseline, as shown in table IV.
Table IV: Comparison with the performance of the baseline on four models across three datasets.
Datasets models ASR(%) BPD(%)
@@ V-C Comparison with baseline
Cora GAE 81.25 99.00 2.82 -1.30
VGAE 83.22 99.13 4.65 0.17
ARGA 92.56 97.53 1.08 -0.71
ARVGA 85.25 88.16 2.49 -0.27
CiteSeer GAE 82.25 99.69 3.13 0.32
VGAE 89.59 99.85 8.51 0.02
ARGA 98.78 99.81 4.97 -0.40
ARVGA 81.55 98.88 1.60 0.37
CS GAE 50.61 98.36 1.13 0.59
VGAE 67.26 98.63 0.95 0.03
ARGA 63.52 97.72 0.32 -0.03
ARVGA 68.67 93.72 2.86 0.30
From the experimental results in table IV, the performance of the proposed backdoor attack against four models is better than that of the baseline on all three datasets, with four backdoor models achieving ASRs of over 88% and BPD fluctuation of less than 1.5% on the three datasets. Taking the Cora dataset as an example, our method achieved 99.00%, 99.13%, 97.53%, and 88.16% ASRs respectively on the four backdoored models of GAE, VGAE, ARGA and ARVG, which are higher than the baseline’s 81.25%, 83.22%, 92.56%, and 85.25%, respectively. All of our results on the Citeseer dataset achieve an ASR of more than 99%. On the CS dataset, the ASRs of baseline drops dramatically, while our method still achieves an over 98% ASRs at CS.
2. Comparison of the performance of the proposed backdoor attack with that of the baseline on GAE using large -scale dataset.
Considering that graph data in the real world are complex, the baseline evaluates its performance on the large-scale dataset physcics. It only tested the performance of the backdoored GAE model with the poisoning rate of 10% on the dataset physics. So we choose from table III the results of backdoored GAE model with a poisoning rate of 1% on the dataset physics to compare with it. The comparison is shown in table V.
From table V, we can see that the performance of the proposed backdoor attack exceeds that of the baseline, with a 60% increase in ASR and a 0.78% reduction in BPD.
Table V: Comparison with the baseline on the large-scale dataset
Method ASR(%) BPD(%)
baseline 39.02 0.78
ours 99.94 -0.44
Table VI: The impact of trigger size on ASR and BPD.
Datasets $\lambda$(%) ASR(%) BPD(%)
@@
Cora 0.2 94.31 96.14 84.01 66.48 -1.27 0.18 -0.69 -0.28
0.5 98.48 98.63 93.52 75.18 -1.29 0.21 -0.70 -0.28
1 99.98 99.13 97.53 88.15 -1.28 0.17 -0.71 -0.27
2 99.44 99.55 98.83 96.42 -1.32 0.20 -0.72 -0.22
5 99.66 99.57 99.46 99.22 -1.38 0.22 -0.72 -0.28
CiteSeer 0.2 99.01 99.13 97.37 69.91 0.76 -0.20 -0.38 0.36
0.5 99.72 99.66 99.49 90.04 0.71 -0.08 -0.35 0.28
1 99.69 99.85 99.81 98.88 0.32 0.02 -0.40 0.37
2 99.95 99.93 99.92 99.63 0.52 0.11 -0.39 0.45
5 99.87 99.92 99.92 99.91 0.68 0.07 -0.39 0.39
CS 0.2 89.96 90.64 84.63 65.42 0.60 0.04 0.02 0.30
0.5 96.91 96.92 94.92 82.51 0.62 0.05 -0.03 0.30
1 98.36 98.63 97.72 93.72 0.59 0.03 -0.03 0.30
2 99.94 99.94 99.92 99.63 0.40 0.09 -0.39 0.45
5 99.87 99.92 99.98 99.63 0.40 0.09 -0.39 0.45
Physics 0.2 97.99 97.33 85.50 68.61 -0.41 -0.04 0.10 -0.59
0.5 99.52 98.95 96.40 89.16 -0.40 -0.04 0.10 -0.59
1 99.94 99.73 98.09 97.10 -0.44 -0.05 0.10 -0.59
2 99.94 99.95 98.76 98.77 -0.47 -0.03 0.11 -0.59
5 99.98 99.99 99.85 99.82 -0.43 -0.05 0.11 -0.59
Figure 4: The impact of trigger size on ASR and BPD.
V-D Testing the impact of different trigger sizes on ASR and BPD
In this section, we evaluate the impact of different trigger sizes on ASR and BPD of the proposed backdoor attack. As described in section V-A4, $k=\lambda d$, where $k$ is the trigger size, which represents the number of the modified elements in the feature vector of the trigger node in the process of generating the trigger node, $d$ is the total number of elements in the trigger node feature vector, and $\lambda$ is the budget, which represents the percentage of the number of modified elements in the feature vector of the trigger node compared to the total number of the elements. We evaluate the impact of different trigger sizes on ASRs and BPDs of four backdoored models of the link predictions models (GAE, VGAE, ARGA and ARVGA) on four datasets (Cora, CiteSeer, CS, and Physics) respectively, and the values of the budget $\lambda$ in each test were 0.2%, 0.5%, 1%, 2%, and 5%, respectively, which corresponds to the change in trigger size. The experimental results are shown in table VI.
Figure 4 is a visualization of the experimental results, where the top four charts show the variation of ASRs with the budgets for the four backdoored models on the four datasets, with the vertical coordinate being ASR and the horizontal coordinate being $\lambda$. The bottom four charts show the variation of BPDs with the budgets for the four backdoored models on the four datasets, with the vertical coordinate being BPD and the horizontal coordinate being $\lambda$ as well. We can see from the figure that the ASRs of four backdoored models(GAE,VGAE,ARGA,ARVGA) on four datasets( Cora, CiteSeer, CS, Physics) increase as $\lambda$ increases. The backdoored models of GAE, VGAE and ARGA achieve a ASR greater than 84% on these four datasets when $\lambda$ is just 0.2%, while the ASR of the backdoored model of ARVGA on these four datasets is greater than 64% and less than 70% when $\lambda$ takes the same value,and the ASR of these four backdoored models on these four datasets achieve ASRs greater than 96% when $\lambda$ is 2%. The above experimental results indicate that the proposed backdoor attack can achieve good attack performance with small trigger size. Besides, we can also see from the figure that the BPDs of these four backdoored models on these four databases fluctuate with the budget $\lambda$, but the fluctuations are less than 1%, which means the trigger size has little impact on the BPDs of these backdoor models and the performance of the backdoored models on clean samples is close to that of corresponding clean models, indicating better stealthiness of the proposed backdoor attack.
Figure 5: The comparison of the ASR of the backoored models obtained by adopting two poisoning methods PBSF and RB.
V-E Ablation Studies
In this section, we will conduct ablation study to confirm whether the method to generate poisoning training dataset through selecting unlinked node pairs with sparse features based on the score function to inject the trigger node(the poisoning method is called Poisoning Based on Score Function for short, PBSF), as described in section 4.3, is crucial to improving the performance of the proposed backdoor attack. For comparison, we firstly replace PBSF with another method to generate poisoning training datasets ,which randomly selects unlinked node pairs to inject the trigger(the poisoning method is called Random Poisoning for short, RB). We poison four training datasets (Cora, CiteSeer, CS and Physics) with the RB method and get the four backdoored models(GAE, VGAE, ARGA and ARVGA) with these poisoning training datasets. Then we test the ASRs of these four backdoored models on four datasets respectively and compare the results with that of the backdoored models obtained by the PBSF method. The poisoning rate is set to 1%. The budget is 0.5%. The experimental results are shown in figure 5, where horizontal coordinates represents four backdoored models obtained by adopting the poisoning methods of PBSF and RB ,the vertical coordinates represent ASRs of these backdoored models, the blue bars and the orange bars in the figure correspond to the ASRs of the backdoored models generated by poisoning methods RB and PBSF respectively. From the figure, we can see that the ASRs of the backdoored models obtained by using PBSF method are higher than those of the backdoored models obtained by using RB method, especially on large dataset CS and Physics, indicating that the poisoning method PBSF is necessary for the proposed backdoor attack to achieve excellent performance. We can also see from the figure that the ASRs of the backdoored models using the two poisoning methods are close to each other on small datasets Cora and CiteSeer. The reason for that is that most of the unlinked node pairs in the small datasets Cora and CiteSeer have sparse features,so most unlinked node pairs randomly selected by RB to inject the trigger node have sparse features with high probability, which achieve similar poisoning to PBSF that selects unlinked node pairs with sparse features by using the score function. Therefore the ASRs of the backdoored models using RB is close to that of the backdoored models using PBSF on small datasets Cora and CiteSeer. In contrast, for large large datasets CS and Physics, where most of the unlinked node pairs do not have sparse features, most unlinked node pairs randomly selected by RB to inject the trigger node are unlikely to have sparse features, and thus the ASRs of the backdoored models using RB is inferior to that of the backdoored models using PBSF.
VI Conclusion
@@
This paper discusses the feasibility of implementing backdoor attacks in link prediction. Specifically, we implement a single-node injection backdoor attack that generate the features of the trigger node based on the differences in feature distribution in the dataset and identifies suitable poisoning node pairs based on the sparsity of feature pairs in the dataset. Our empirical evaluations on four representative link prediction models (GAE, VGAE, ARGA, and ARVGA) and four benchmark datasets demonstrate that our backdoor attack can achieve high ASRs with low poisoning rates, and has minimal impact on the accuracy of clean test graphs. Currently, there is limited research on defenses against backdoor attacks in graphs, and in the future, we will explore defense methods against backdoor attacks in link prediction.
Acknowledgement
The research of the paper was supported by Natural Science Foundation of Shanghai Municipality (Grant NO.22ZR1422600)